A wooden boat is found buried beneath the ocean floor. Scientists believe that timbers making up the boat had 8 g of carbon-14 in them when alive. They now have 2.3 g of carbon-14. How old is the boat if the half-life of carbon-14 is 5715 years?
1
Expert's answer
2020-05-04T12:40:01-0400

Determine the decimal amount remaining: "2.3\/8=0.2875".

Determine the number of half-lives: "(0.5)^n=0.2875",

"n \ln 0.5=\ln 0.2875,\,\,n=\frac{\ln 0.2875}{\ln 0.5}=1.7984".

Determine years elapsed: "5715 \cdot n=10277.66" years.